OpenStudy (anonymous):

Suppose you choose a marble from a bag containing 4 red marbles, 2 white marbles, and 3 blue marbles. You return the first marble to the bag and then choose again. Find P(red and blue). A. 4/3 B. 2/3 C. 7/9 D. 4/27

OpenStudy (mertsj):

\[\frac{4}{9}\times\frac{3}{9}=\frac{4}{27}\]

OpenStudy (ganpat):

i guess, it should be 7/9.. Probability of getting white marble = 2/9.. probability of getting red and blue marble = 1- 2/9...

OpenStudy (mertsj):

Stop being ridiculous.

OpenStudy (ganpat):

excuse me..